# AttentionRAG: Attention-Guided Context Pruning in Retrieval-Augmented Generation

### v1

@@ Abstract
Yixiong Fang1, Tianran Sun1, Yuling Shi1, Xiaodong Gu1 1Shanghai Jiao Tong University Correspondence: xiaodong.gu@sjtu.edu.cn
@@ 1 Introduction
Figure 2: Illustration of AttentionRAG
@@ 3.1 Construct Answer Hint Prefix
We prompt the LLM to construct the answer hint prefix [^1] We use GPT-4o Mini for the generation, the prompt detail can be referred to  §B.. In detail, according to the query’s grammatical attributes, we categorize the answer hint prefix into two types : empty and non-empty ones. For example, queries like wh-questions, the hint prefix can be derived by the query itself. In contrast, queries like yes/no-questions, where the answer’s first token is "Yes/No", already align with the next-token-prediction paradigm. Leveraging the semantic understanding capabilities of LLMs, we prompt them with example answer hint prefix of various types questions to automatically determine the type and generate the answer prefix hint.
@@ 3.2 Compute Attention Features from LLM
In this part, we aim to compute attention features based on the focal token. To address the issue of diluted attention scores, we divide the retrieved context $C$ produced by the RAG framework into smaller chunks. We adopt a uniform chunking strategy, assuming each chunk consists of $m$ tokens. Let $c_{1},\dots,c_{n}$ represent the resulting chunks, where $n=\lceil|C|/m\rceil$. For each chunk $c_{j}$, we concatenate the chunked context, query, instruction(we instruct the LLM to generate "none" after hint prefix if the chunk is irrelevant, which will be used in §3.3) and the generated answer hint prefix and feed this into the LLM. The LLM is then instructed to perform next-token prediction and compute the attention scores.
@@ 3.4 Time Efficiency and Batch Generation
Since we employ a next-token prediction paradigm, only one focal token needs to be generated for each chunk. Furthermore, as each chunk is processed independently, batch generation can be used to accelerate the process. This approach results in high time efficiency.
We provide the pseudocode of our method in Algorithm 1 in Appendix.
@@ 4.1 Datasets
Due to fluctuations in the experimental results, each experiment was conducted three times, and the final result is the average of these trials. We evaluate AttentionRAG in two key aspects: long context understanding and long context reasoning, reflected in two suites of benchmarks respectively:
LongBench
(Bai et al., 2024): LongBench provides an extensive framework to evaluate an LLM’s ability to comprehend and process long-form texts. It encompasses a diverse suite of tasks—including complex information extraction, multi-step reasoning, and coherent text generation—that are designed to challenge models when key details are interspersed throughout extended passages. For our experiments, we specifically incorporate datasets such as TriviaQA (Joshi et al., 2017), HotpotQA (Yang et al., 2018), and 2WikiMQA (Ho et al., 2020), as each brings unique demands: TriviaQA assesses fact retrieval over long contexts, HotpotQA emphasizes multi-hop reasoning across dispersed clues, and 2WikiMQA tests the model’s ability to synthesize information from multiple sources. To ensure zero-shot consistency in testing, we do not use the example section from LongBench/TriviaQA.
@@
(Kuratov et al., 2024): The BABILong benchmark provides a comprehensive framework for evaluating an LLM’s ability to reason over long contexts. It includes 20 varied tasks, such as fact chaining, simple induction, deduction, counting, and list/set manipulation, which become increasingly challenging when the relevant information is spread across extended natural texts. To assess AttentionRAG’s reasoning capabilities, we select test splits of 1kqa1, 2kqa1, and 4kqa1 tokens to demonstrate its performance across different context lengths.
@@ 4.2 Metrics
LLM-as-a-judge scores: We leverage GPT-4 (et al., 2024) to assess the correctness of the model-generated answers. Specifically, we input the question, the model-generated answer, and the ground truth answer into GPT-4, asking it to determine whether the provided answer is correct. The prompt used for the scoring can be found in  §B.
Compression Ratio (CR): The ratio of tokens in the original context to the compressed context, defined as:
$CR=\frac{\text{\# tokens in original context}}{\text{\# tokens in compressed context}}$
@@ 4.3 Baselines
We compare AttentionRAG with the LLMLingua family, which includes LLMLingua (Jiang et al., 2023), LongLLMLingua (Jiang et al., 2024), and LLMLingua2 (Pan et al., 2024). Specifically, we choose LLMLingua2 and LongLLMLingua for baseline. They are state-of-the-art context compression methods. These baselines consist of question-aware methods (LongLLMLingua) and question-unaware methods (LLMLingua and LLMLingua2). The variety in baseline selection highlights the robustness of AttentionRAG. To ensure a fair comparison, we maintain a compression rate for the LLMLingua family that is equal to or greater than that of AttentionRAG. Additionally, we compare our results against the uncompressed context for reference, denoted as “Original Prompt”, and random compression, denoted as “Random”.
@@ 5.1 Overall Results
Model Method 2WikiMQA HotpotQA TriviaQA
@@
Llama-3.1-8B-Instruct Original Prompt 0.49 0.46 1.0x 0.50 0.62 1.0x 0.87 0.80 1.0x
@@
LLMLingua2 (small) 0.24 0.26 5.0x 0.40 0.45 3.3x 0.86 0.82 1.7x
LLMLingua2 (large) 0.33 0.30 5.0x 0.44 0.51 3.3x 0.84 0.76 1.7x
LongLLMLingua 0.23 0.22 3.3x 0.33 0.40 3.3x 0.83 0.72 1.7x
AttentionRAG (small) 0.39 0.36 6.3x 0.45 0.55 3.7x 0.84 0.77 2.0x
AttentionRAG (large) 0.42 0.38 15x 0.48 0.61 5.6x 0.89 0.81 1.7x
Qwen2.5-7B-Instruct Original Prompt 0.56 0.46 1.0x 0.58 0.68 1.0x 0.94 0.82 1.0x
@@
LLMLingua2 (small) 0.29 0.27 5.0x 0.43 0.47 3.3x 0.91 0.80 1.7x
LLMLingua2 (large) 0.40 0.28 5.0x 0.48 0.58 3.3x 0.87 0.81 1.7x
LongLLMLingua 0.31 0.18 4.0x 0.28 0.31 3.3x 0.83 0.73 1.7x
AttentionRAG (small) 0.41 0.28 6.3x 0.51 0.54 3.7x 0.83 0.73 2.0x
AttentionRAG (large) 0.41 0.30 15x 0.53 0.56 5.6x 0.88 0.80 1.7x
Table 1: Performance of different methods in 2WikiMQA, HotpotQA and TriviaQA dataset.
@@
As the results in Tables 1 and 2 show, AttentionRAG outperforms nearly all baseline methods across the benchmarks, consistently achieving superior results. On the LongBench benchmark, which includes contexts of tens of thousands of tokens, AttentionRAG outperforms the LLMLingua methods in most metrics, particularly in BABILong 1k, where it achieves an 18% higher score than the best LLMLingua model. Specifically, in TriviaQA, AttentionRAG outperforms the uncompressed method by 2% in EM score, demonstrating the effectiveness of our approach. However, the low compression ratio is due to two parallel factors: the scattering of relevant information across the context and the use of uniform hyperparameters. All these factors contribute to the lower ratio, which we aim to enhance. We discuss this in more detail in the Ablation Study.
Similar results can be observed in the BABILong benchmark, which involves a wider range of RAG context lengths. Our method consistently outperforms all LLMLingua methods while achieving the highest compression ratio, further demonstrating the effectiveness of AttentionRAG in varying context sizes.
@@ 5.2 Case Study
LLMLingua2: Waldrada Lotharingia mistress Lothair II Gallo sister Thietgaud niece Gunther Vita Sancti related Eberhard II Etichonids 855 Lothar II married Teutberga 858 862 Nicholas 863Charles …(632 tokens) Munkács.$\times$
@@ 5.3 Ablation Study
Efficiency Analysis
As discussed in  §3, we divide the context into smaller chunks and use attention mechanisms for compression. To speed up the process, we can design the prompt with an answer template that allows the overall generation time to be equal to the forward pass time of the compression model plus the compressed generation time of the generation model. Since a lightweight model is used for compression, the overall generation time is significantly reduced.
Attention Layer Choice In LLMs, attention mechanisms across different layers capture varying levels of information. Specifically, earlier layers often focus on syntactic structures, while deeper layers tend to capture more abstract semantic representations (Ben-Artzy and Schwartz, 2024). This hierarchical processing enables LLMs to effectively understand and generate complex language patterns (Jin et al., 2025). We adopt a comprehensive approach by aggregating attention scores across all layers for compression. As discussed in  §3.2, this method ensures that the compression process leverages the full spectrum of information captured by the model and mitigate the potential attention bias across layers. We compare our aggregated approach using all layers with methods using different layer subsets (shallow, middle, and deep). To evaluate the effectiveness of these approaches, we conduct experiments on the HotpotQA dataset using the Llama-3.1-8B-Instruct model. As shown in Table 4, our method of summing attention scores across all layers outperforms the subset-layer approaches, supporting our hypothesis that leveraging the full spectrum of information captured across all layers provides better performance.
Hyperparameters We conduct ablation studies on two hyperparameters: the chunk size that we further partition the retrieved lengthy contexts (§3) and the number of Tok-$K$ tokens for selecting the sentence in each chunk in the attention-based compression process (§3.3). For longer contexts (e.g., LongBench tasks), we increase the chunk size and use higher top-K values, as the information is more dispersed and larger chunks require a higher top-K to avoid internal bias. In contrast, for shorter contexts (e.g., Babilong tasks), we use smaller chunk sizes and lower top-K values to better capture the relevant information.  [^2] The hyperparameter settings for these experiments are detailed in  §B. We conducted experiments on the TriviaQA dataset, separately testing various configurations of chunk sizes and K values. As shown in Table 5, we find that decreasing K and increasing the chunk size significantly reduce the compression ratio, while only slightly affecting the model’s performance. This is because these factors may decrease the relative context retrieved in each chunk. Therefore, in addition to a fixed compression ratio, we introduce more dynamic parameters to better control the balance between efficiency and performance.
@@
0 - 10 0.35 0.43 4.5x
11 - 20 0.38 0.50 3.6x
21 - 31 0.40 0.48 3.7x
0 - 31 0.42 0.54 3.6x
Table 4: Performance on HotpotQA with different subset of layers
@@
Size of Compression Models To assess the effectiveness of our lightweight compression model, we compare the results of compressing LLaMA-3.1-Instruct 8B against LLaMA-3.1-Instruct 70B. Our findings indicate that within the attention mechanism, the lightweight model delivers performance comparable to the larger model, suggesting that smaller models can achieve similar results. As shown in Table 1, the difference in performance between the large and small models is small, highlighting that when only attention scores are required, the smaller model is sufficiently capable for the compression task and can significantly enhance efficiency.
@@ 6 Related Work
RAG has demonstrated effectiveness in various domains, such as open-domain question-answering (Han et al., 2024). However, the presence of redundant and irrelevant retrieved information can significantly hinder the performance of LLMs (Shi et al., 2023). To address this challenge, numerous methods have been proposed to improve the quality of retrieved information Wang et al. (2023); Xu et al. (2023). For instance, Wang et al. (2023) presents a technique that trains models to identify and filter out irrelevant contexts in RAG systems, thereby enhancing both retrieval accuracy and downstream task performance. Xu et al. (2023), on the other hand, proposes a method that leverages context compression by training extraction models, which retain critical information while minimally impacting performance. In contrast to these approaches, AttentionRAG relies on the inherent ability of LLMs to recognize useful information within their hidden states without requiring additional training, and can be applied to common compression problems. Our method not only achieves superior performance but is also more transferable and readily applicable across different models.
@@
To address the efficiency and cost challenges of long-context generation, various prompt compression methods have been proposed. These methods can be broadly categorized into soft prompts and hard prompts. In the domain of soft prompts, Mu et al. (2024) introduces the gist token, achieving a high compression rate with minimal performance loss. Similarly, Li et al. (2024) compresses long contexts into a single special token, retaining 50% to 70% of the original performance while significantly reducing input length. For hard prompts, Jiang et al. (2023) presents LLMLingua, a method that identifies and removes unimportant tokens from prompts, optimizing inference efficiency. Building upon this, Jiang et al. (2024) proposes LongLLMLingua, which further enhances the ability of language models to capture key information in long-context scenarios through prompt compression. This approach achieves up to a 17.1% performance improvement with a 4x compression ratio, effectively addressing the computational and latency challenges of processing long contexts. Additionally, Pan et al. (2024) introduces LLMLingua-2, a task-agnostic prompt compression method trained via data distillation from GPT-4 for token classification with a BERT-level encoder. LLMLingua-2 excels in handling out-of-domain data, offering 3x-6x faster performance compared to its predecessors. In contrast to these methods, AttentionRAG focuses on improving LLM performance by leveraging explainable contexts derived from internal features. Our experiments will compare AttentionRAG with LLMLingua2 and LongLLMLingua to highlight its effectiveness and efficiency.

### v2

@@ Abstract
Yixiong Fang1thanks:   Equal contribution., Tianran Sun1111Corresponding author., Yuling Shi1, Xiaodong Gu1thanks:   Corresponding author. Shanghai Jiao Tong University {fangyixiong, Seriousss, yuling.shi, xiaodong.gu}@sjtu.edu.cn
@@ 1 Introduction
Figure 2: Illustration of AttentionRAG: 1) We first generate a hint prefix using the original query; if generation fails, we fallback to a fixed hint prefix. 2) The long context is then divided into smaller chunks for localized processing. 3) For each chunk, we batch the input with the hint prefix (1 token only) and use an LLM to compute attention features over the full context, focusing on the generated anchor token. 4) Using these attention features, we identify and extract the most relevant content from each chunk to construct a compressed context.
@@ 3.1 Construct Answer Hint Prefix
We prompt the LLM to construct the answer hint prefix [^1] We use GPT-4o Mini for the generation, the prompt detail and fixed hint prefix can be referred to  §B.1 and  §B.2.. In detail, according to the query’s grammatical attributes, we categorize the answer hint prefix into two types: empty and non-empty ones. For example, for queries like wh-questions, the hint prefix can be derived from the query itself. In contrast, queries like yes/no-questions, where the answer’s first token is "Yes/No", already align with the next-token-prediction paradigm. However, some type of questions may not have such hint prefix (like summarization task), we also provide a fixed hint prefix that prompt the model to output the keyword of context, ensuring robustness of our method. Our fixed hint prefix is: "Please output the most relevant keyword or phrase that is relevant to the answer of the question." Leveraging the semantic understanding capabilities of LLMs, we prompt them with example answer hint prefix of various types questions to automatically determine the type and generate the answer prefix hint.
@@ 3.2 Compute Attention Features from LLM
In this part, we aim to compute attention features based on the focal token. To address the issue of diluted attention scores, we divide the retrieved context $C$ produced by the RAG framework into smaller chunks. We adopt a uniform chunking strategy, assuming each chunk consists of $m$ tokens. Let $c_{1},\dots,c_{n}$ represent the resulting chunks, where $n=\lceil|C|/m\rceil$. For each chunk $c_{j}$, we concatenate the chunked context, query, and instruction (we instruct the LLM to generate "none" after the hint prefix if the chunk is irrelevant, which will be used in §3.3) and the generated answer hint prefix and feed this into the LLM. The LLM is then instructed to perform next-token prediction and compute the attention scores.
@@ 3.4 Time Efficiency and Batch Generation
Since we employ a next-token prediction paradigm, only one focal token needs to be generated for each chunk. Furthermore, as each chunk is processed independently, batch generation can be used to accelerate the process. This approach results in high time efficiency. Moreover, we can use quantified model to further accelerate the compression process. We provide the pseudocode of our method in Algorithm 1 in Appendix.
@@ 4.1 Datasets
Due to fluctuations in the experimental results, each experiment was conducted three times, and the final result is the average of these trials.
Long Context Reasoning
For our experiments, we incorporate datasets TriviaQA (Joshi et al., 2017), HotpotQA (Yang et al., 2018), and 2WikiMQA (Ho et al., 2020) from Longbench (Bai et al., 2024): TriviaQA assesses fact retrieval over long contexts, HotpotQA emphasizes multi-hop reasoning across dispersed clues, and 2WikiMQA tests the model’s ability to synthesize information from multiple sources.
@@
(Kuratov et al., 2024): The BABILong benchmark provides a comprehensive framework for evaluating an LLM’s ability to reason and retrieve over long contexts. To assess AttentionRAG’s capabilities on relatively shorter and sparse content, we select test splits of 1kqa1, 2kqa1, and 4kqa1 tokens to demonstrate its performance across different circumstances.
Long Context Summarization
To demonstrate the robustness of AttentionRAG, we also evaluate it on GovReport (Huang et al., 2021) from Longbench. In this task, we employ fixed hint prefix, showing the stability of our method.
@@ 4.2 Metrics
LLM-as-a-judge scores: We leverage GPT-4 et al. (2024) to assess the correctness of the model-generated answers. Specifically, we input the question, the model-generated answer, and the ground truth answer into GPT-4o, asking it to determine whether the provided answer is correct. The prompt used for the scoring can be found in  §B.
[…]
@@ 4.3 Baselines
We compare AttentionRAG with state-of-the-art context compression baselines, including LLMLingua2 (Pan et al., 2024) and LongLLMLingua (Jiang et al., 2024), which represent question-unaware and question-aware approaches, respectively. This diverse selection underscores the robustness of AttentionRAG. For fair comparison, we match or exceed their compression rates. We also include the uncompressed context (“Original Prompt”), a random compression baseline (“Random”), and a vanilla RAG baseline using BGE-M3 (Chen et al., 2024).
@@ 5.1 Overall Results
Model Method 2WikiMQA HotpotQA TriviaQA Gov
@@
Llama-3.1-8B-Instruct Original Prompt 0.49 0.46 1.0x 0.50 0.62 1.0x 0.87 0.80 1.0x 0.32 1.0x
@@
LLMLingua2 (small) 0.24 0.26 5.0x 0.40 0.45 3.3x 0.86 0.82 1.7x 0.23 3.3x
LLMLingua2 (large) 0.33 0.30 5.0x 0.44 0.51 3.3x 0.84 0.76 1.7x 0.24 3.3x
LongLLMLingua 0.23 0.22 3.3x 0.33 0.40 3.3x 0.83 0.72 1.7x 0.23 3.1x
BGE-M3 0.40 0.35 3.5x 0.42 0.50 3.0x 0.80 0.77 1.9x 0.21 3.0x
AttentionRAG (small) 0.39 0.36 6.3x 0.45 0.55 3.7x 0.84 0.77 2.0x 0.28 3.4x
AttentionRAG (large) 0.42 0.38 15x 0.48 0.61 5.6x 0.89 0.81 1.7x 0.28 4.2x
Qwen2.5-7B-Instruct Original Prompt 0.56 0.46 1.0x 0.58 0.68 1.0x 0.94 0.82 1.0x 0.31 1.0x
@@
LLMLingua2 (small) 0.29 0.27 5.0x 0.43 0.47 3.3x 0.91 0.80 1.7x 0.21 3.3x
LLMLingua2 (large) 0.40 0.28 5.0x 0.48 0.58 3.3x 0.87 0.81 1.7x 0.22 3.3x
LongLLMLingua 0.31 0.18 4.0x 0.28 0.31 3.3x 0.83 0.73 1.7x 0.23 3.1x
BGE-M3 0.40 0.35 3.5x 0.42 0.50 3.0x 0.80 0.77 1.9x 0.22 3.0x
AttentionRAG (small) 0.41 0.28 6.3x 0.51 0.54 3.7x 0.83 0.73 2.0x 0.28 3.4x
AttentionRAG (large) 0.41 0.30 15x 0.53 0.56 5.6x 0.88 0.80 1.7x 0.28 4.2x
Table 1: Performance comparison on QA datasets (2WikiMQA, HotpotQA, TriviaQA) and the Gov summarization dataset.
@@
As the results in Tables 1 and 2 show, AttentionRAG outperforms nearly all baseline methods across the benchmarks, consistently achieving superior results. On the LongBench benchmark, which includes contexts of tens of thousands of tokens, AttentionRAG outperforms the LLMLingua methods in most metrics, particularly in BABILong 1k, where it achieves an 18% higher score than the best LLMLingua model. Specifically, in TriviaQA, AttentionRAG outperforms the uncompressed method by 2% in EM score, demonstrating the effectiveness of our approach. However, the low compression ratio is due to two parallel factors: the scattering of relevant information across the context and the use of uniform hyperparameters. All these factors contribute to the lower ratio, which we aim to enhance. We discuss this in more detail in the Ablation Study. Also, our method excels in summarization tasks, showing the stability of our method. Similar results can be observed in the BABILong benchmark, which involves a wider range of RAG context lengths. Our method consistently outperforms all LLMLingua methods while achieving the highest compression ratio, further demonstrating the effectiveness of AttentionRAG in varying context sizes.
5.2 Efficiency Analysis
As discussed in §3, we divide the context into smaller chunks and use attention mechanisms for compression. The overall time cost is spent in the forward pass of the compression model plus the answering of the generation model. We employ LLaMA-3.1-Instruct with int4 quantization for compression. We choose HotpotQA, where the average context length is about tens of thousands tokens. As shown in Table 3, the overall time cost is significantly reduced compared to baseline methods, and the quantized model can also ensure similar performance on results.
Method EM $\uparrow$ LLM Judge $\uparrow$ CR $\uparrow$ Latency
Compression Answering
Original Prompt 0.50 0.62 1.0x 19.36
Vanilla RAG (BGE-M3) 0.42 0.50 3.0x 6.11 5.99
LLMLingua2 (small) 0.40 0.45 3.3x 0.44 5.39
LLMLingua2 (large) 0.44 0.51 3.3x 0.88 5.43
LongLLMLingua 0.42 0.52 3.3x 7.12 5.41
AttentionRAG 0.45 0.54 4.0x 3.99 4.94
Table 3: Time latency for inference by various methods on HotpotQA
@@ 5.3 Case Study
[…]
@@ 5.4 Ablation Study
Fixed Hint Prefix As discussed in §3.1, we use a fixed hint prefix for questions that cannot generate one dynamically. The summarization performance under this setting is reported in Table 1. To further validate the robustness of this approach, we conduct experiments on three additional datasets: 2WikiMQA, HotpotQA, and TriviaQA. As shown in Table 5, we experiment with LLaMA-3.1-8B-Instruct, using only a fixed hint prefix results in a slight performance drop compared to our original method, yet still consistently outperforms all baselines. This demonstrates that even a static hint prefix serves as an effective anchor token, reliably guiding the model to identify important information within the context.
@@
HotpotQA 0.44 0.53 3.3x
2WikiMQA 0.44 0.37 5.2x
TriviaQA 0.81 0.75 2.0x
Table 5: Performance of AttentionRAG using fixed hint prefix across datasets.
Combination with other RAG methods To explore the integration potential of our method with retrieval-based techniques, we implemented a two-stage pipeline: we first use vanilla RAG (BGE-M3) for initial context retrieval, followed by our method for compression. Specifically, we retain the RAG chunk size at 300 tokens, while retrieve the top 15 relevant chunks, and then apply our compression method. We use LLaMA-3.1-8B-Instruct as the generation model and conduct experiments on the HotpotQA, 2WikiMQA, and TriviaQA benchmarks (See Table 6).
EM LLM Judge CR
HotpotQA 0.46 0.58 9.1x (+5.4x)
2wikimqa 0.43 0.40 7.9x (+1.6x)
triviaqa 0.88 0.83 3.2x (+1.5x)
Table 6: Combination with other RAG method
Hyperparameters We conduct ablation studies on two hyperparameters: chunk size (§3) and the number of Top-$K$ tokens used to select sentences within each chunk (§3.3). For long contexts (e.g., LongBench), we use larger chunks and higher $K$ to handle dispersed information; for shorter contexts (e.g., Babilong), we use smaller chunks and lower $K$ for finer granularity. [^2] Details in §B. Using TriviaQA, we test various configurations and observe (Table 7) that increasing chunk size and reducing $K$ lowers the compression ratio with minimal performance drop. This suggests that dynamic hyperparameter tuning offers a better trade-off between efficiency and accuracy than fixed-ratio compression.
@@
Size of Compression Models To evaluate the model-size sensitivity of our compression approach, we compare the performance of AttentionRAG when using LLaMA-3.1-Instruct 8B versus 70B to compute attention scores. As shown in Table 1, the performance gap is minimal, indicating that even a lightweight model can effectively capture the attention patterns needed for compression. This suggests that AttentionRAG is largely agnostic to model size and can maintain strong performance without relying on large-scale models—offering substantial efficiency benefits. Furthermore, as demonstrated in §5.2, our method can be further accelerated through quantization. Overall, AttentionRAG is highly scalable and future-proof, with the potential to continuously benefit from advances in foundation model development while remaining efficient and adaptable.
@@ 6 Related Work
RAG has shown strong performance in tasks like open-domain QA (Han et al., 2024; Liu et al., 2024), visual QA (Wang et al., 2025a, b), and many other tasks (Qiu et al., 2025). But its effectiveness is often hindered by noisy or redundant retrieved content (Shi et al., 2023). To address this, recent work has focused on improving retrieval quality. Wang et al. (2023) trains models to filter irrelevant content, while Xu et al. (2023) uses extraction-based compression to retain key information. Unlike these approaches, AttentionRAG does not require additional training. It leverages internal attention signals to identify informative content, offering strong performance and broad transferability across tasks and models.
@@
To reduce the cost of long-context generation, both soft and hard prompt compression methods have been proposed. Soft prompt methods include Gist (Mu et al., 2024) and 500x Compressor (Li et al., 2024), which compress context into dense tokens with minimal loss. Hard prompt approaches like LLMLingua (Jiang et al., 2023), LongLLMLingua (Jiang et al., 2024), and LLMLingua2 (Pan et al., 2024) use token-level filtering, achieving substantial speedups and compression. In contrast, AttentionRAG enhances LLM performance by selecting explainable, model-attended content without retraining. We compare against LongLLMLingua and LLMLingua2 to demonstrate its efficiency and robustness.